OpenStudy (anonymous):

Part 1: Describe the graphs of the functions f(x) = 2x – 1 and g(x) = –2x – 1. (3 points) Part 2: Compare and contrast the domain and range of f(x) and g(x). (3 points) help please guysss

OpenStudy (turingtest):

\[f(x)=2x-1\\g(x)=-2x-1\]?

OpenStudy (anonymous):

i dont get how to do this can help me please?

OpenStudy (anonymous):

pleaseeeee i really need thisss for a finall gradee!

OpenStudy (turingtest):

I'm just asking if I got the problem right do you know about equations like\[y=mx+b\]? those equations are called "linear", for one reason because they represent a line in space

OpenStudy (turingtest):

is that formula above completely new to you?

OpenStudy (anonymous):

no i know which one it is i just dont the problem , and i forgot the steps

OpenStudy (turingtest):

since all they want is a description you just need to recognize the parts the equations you have are of the form y=mx+b what does m represent, do you remeber?

OpenStudy (turingtest):

remember*

OpenStudy (anonymous):

no i dont remember anything can you do the problem step by step please? i cant remember

OpenStudy (turingtest):

there aren't really steps here, it's an ability to recognize a situation

OpenStudy (turingtest):

not all math problems are just about remembering a procedure in this case you need to remember the "slope-intercept form" y=mx+b and realize "oh, that looks like my equation, so it must be a line" than start thinking about the parts of that line

OpenStudy (turingtest):

you have two equations of this form, so that represents two lines m represents the slope of the line in the formula y=mx+b so what is the difference between the slopes of the two lines ?

OpenStudy (anonymous):

so what is the answer? my text is timed...

OpenStudy (anonymous):

test*

OpenStudy (turingtest):

that is not how this site works I'm afraid we don't just provide answers, and we often don't help on tests. certainly if your class has a provision about using a site like ours while taking a timed test that is cheating, and is not encouraged here

OpenStudy (turingtest):

if you need help in a subject we are happy to provide it, but do not expect a timed result we are about teaching. please come to us in advance so you are not under such pressure when we are trying to help you

OpenStudy (anonymous):

alright thanks .i just dont geet how to do the problem and all i needed was to get how to do it or what it was asking. thanks anyway !

OpenStudy (turingtest):

I am explaining find m that is the slope of each line the question is asking "how are they different?" the next question is asking you if there are any values of x that make y undefined i.e. are there any x that if you plug in to the formula you get something/0, square root of a negative number, or the logarithm of a number less-than or equal-to zero

OpenStudy (turingtest):

if there are no x's you could put in that would cause one of the problems, then the function is defined for all reals if not, then you have to tell what numbers for x are "not allowed" (in the sense I described above)